# Finite basis Kolmogorov-Arnold networks: domain decomposition for data-driven and physics-informed problems

### v1

@@ 5 Conclusions
We have developed domain decomposition-based KAN models for data-driven and physics-informed training with KANs; in accordance with FBPINNs, we denote them as FBKANs. FBKANs are scalable to complex problems and have a strong advantage over other domain decomposition-based approaches in that they do not require enforcement of transmission conditions between the subdomains via the loss function. They allow accurate training using an ensemble of small KANs combined using partition of unity functions, instead of a single large network.
One advantage of FBKANs is that they can be combined with existing techniques to improve the training of KANs and PI-KANs, including residual-based attention weights as introduced in [7], cKANs [13], deep operator KANs [6], and others. These methods would be directly compatible with the FBKAN framework. In future work, we will further examine the scalability of FBKANs and consider their application to higher dimensional problems. We will also consider multilevel FBKANs, following the multilevel FBPINNs [43] approach. Multilevel FBPINNs show improvement over FBPINNs for a large number of subdomains by providing a mechanism for global communication between subdomains. Multilevel FBKANs could offer a similar advantage, allowing robust training with an increasing number of subdomains.
@@ 6 Code and data availability
All code, trained models, and data required to replicate the examples presented in this paper will be released upon publication. Meanwhile, we have released code and Google Colab tutorials for FBKANs in Neuromancer [67] at https://github.com/pnnl/neuromancer/tree/feature/fbkans/examples/KANs, for the reader to explore the ideas implemented in this work.
7 Acknowledgements
The KAN diagrams in fig. 1 were developed with pykan [4]. This project was completed with support from the U.S. Department of Energy, Advanced Scientific Computing Research program, under the Scalable, Efficient and Accelerated Causal Reasoning Operators, Graphs and Spikes for Earth and Embedded Systems (SEA-CROGS) project (Project No. 80278) and under the Uncertainty Quantification for Multifidelity Operator Learning (MOLUcQ) project (Project No. 81739). The computational work was performed using PNNL Institutional Computing at Pacific Northwest National Laboratory. Pacific Northwest National Laboratory (PNNL) is a multi-program national laboratory operated for the U.S. Department of Energy (DOE) by Battelle Memorial Institute under Contract No. DE-AC05-76RL01830.
@@ Appendix A Training parameters
All results in this paper are implemented in JAX [65] using the Jax-KAN [66] KAN implementation. All networks are trained with the ADAM optimizer. For all FBKANs we take the domain overlap $\delta=1.9$.
A.1 Test 1
parameter section 3.1.1 section 3.1.2
@@ A.1 Test 1
Table 3: Hyperparameters used for the results in section 3.1.
A.2 Test 2
@@
Iterations 2400 2400
$N_{data}$ 10000 10000
Table 4: Hyperparameters used for the results in section 3.2. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update. KAN-1/FBKAN-1 use fixed grids.
A.3 Test 3
@@
Table 5: Hyperparameters used for the results in section 4.1. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update.
A.4 Test 4
@@
$\lambda_{ic}$ 1 1 1
Table 6: Hyperparameters used for the results in section 4.2. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update.
A.5 Test 5
@@
Table 7: Hyperparameters used for the results in section 4.3.

### v2

@@ 6 Conclusions
We have developed domain decomposition-based architectures for data-driven and physics-informed training with KANs; in accordance with finite basis physics-informed neural networks (FBPINNs), we denote them as finite basis Kolmogorov–Arnold networks (FBKANs).The finite basis approach enables scalability for complex problems, and they have a strong advantage over other domain decomposition-based approaches in that they do not require enforcement of transmission conditions between the subdomains via the loss function. They allow accurate training using an ensemble of small KANs combined using partition of unity functions, instead of a single large KAN model. We also considered MLFBKANs, following the multilevel FBPINNs [dolean_multilevel_2024] approach. Compared to the one-level approach, MLFBKANs show improvements for training with larger numbers of subdomains, which is particularly interesting for high or multi frequency problems.
Furthermore, FBKANs can be combined with existing techniques to improve the training of KANs and physics-informed KANS (PI-KANs), including residual-based attention weights as introduced in [shukla2024comprehensive], cKANs [ss2024chebyshev], deep operator KANs [abueidda2025deepokan], and others. In future work, we will further examine the scalability of FBKANs and consider their application to higher dimensional problems.
@@ 7 Code and data availability
All code and data required to replicate the examples presented in this paper will be released upon publication. Meanwhile, we have released code and Google Colab tutorials for FBKANs in Neuromancer [Neuromancer2023] at https://github.com/pnnl/neuromancer/tree/feature/fbkans/examples/KANs for the reader to explore the ideas implemented in this work.
8 Acknowledgments
The KAN diagrams in fig.˜1 were developed with pykan [liu2024kan]. This project was completed with support from the U.S. Department of Energy, Advanced Scientific Computing Research program, under the Scalable, Efficient and Accelerated Causal Reasoning Operators, Graphs and Spikes for Earth and Embedded Systems (SEA-CROGS) project (Project No. 80278) and under the Uncertainty Quantification for Multifidelity Operator Learning (MOLUcQ) project (Project No. 81739). The computational work was performed using PNNL Institutional Computing at Pacific Northwest National Laboratory. Pacific Northwest National Laboratory (PNNL) is a multi-program national laboratory operated for the U.S. Department of Energy (DOE) by Battelle Memorial Institute under Contract No. DE-AC05-76RL01830.
@@ Appendix A Training parameters
All results in this paper are implemented in JAX [jax2018github] using the Jax-KAN [Rigas_jaxKAN_A_JAX-based_2024] KAN implementation. All networks are trained with the ADAM optimizer. For all FBKANs we take the domain overlap $\delta=1.9$, based on the values used in [dolean_multilevel_2024].
A.1 Data-driven test 1
parameter section˜3.1.1 section˜3.1.2
@@ A.1 Data-driven test 1
Table 4: Hyperparameters used for the results of the data-driven test 1 in section˜3.1.
A.2 Data-driven test 2
@@
# iterations $2400$ $2400$
$N_{data}$ 10 00010\,000 10 00010\,000
Table 5: Hyperparameters used for the results of the data-driven test 2 in section˜3.2. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update. KAN-1 / FBKAN-1 use fixed grids.
A.3 Physics-informed test 1
@@
Table 6: Hyperparameters used for the results of the physics-informed test 1 in section˜4.1. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update.
A.4 Physics-informed test 2
@@
$\lambda_{bc}$ 1 1 1
Table 7: Hyperparameters used for the results of the physics-informed test 2 in section˜4.2. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update.
A.5 Physics-informed test 3
@@
Table 8: Hyperparameters used for the results of the physics-informed test 3 in section˜4.3.
A.6 Multilevel physics-informed test 1
parameter
KAN architecture [2, 10, 1]
$L$ 4 – 36
NN 1 – 4
$g$ 5
$g$ schedule -
$k$ 5
initial learning rate 0.005
learning rate scale -
# iterations 30 00030\,000
$N_{r}$ 800
$N_{bc}$ 400
$\lambda_{r}$ 0.01
$\lambda_{bc}$ 1
Table 9: Hyperparameters used for the results of the multilevel physics-informed test 1 in section˜5.1. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update.
A.7 Multilevel physics-informed test 2
parameter
KAN architecture [2, 10, 1]
$L$ 4 – 36
NN 1 – 4
$g$ 5
$g$ schedule -
$k$ 5
initial learning rate 0.005
learning rate scale -
# iterations 30 00030\,000
$N_{r}$ 800
$N_{bc}$ 400
$\lambda_{r}$ 0.001
$\lambda_{bc}$ 1
Table 10: Hyperparameters used for the results of the multilevel physics-informed test 2 in section˜5.2. The grid ($g$) schedule denotes the iterations at which the grid is updated. The learning rate scale denotes the change to the learning rate at each grid update.